OpenStudy (littlebird):

Natural silver ore contains only two different isotopes, silver-­‐107 and silver-­‐109. How many atoms of silver-­‐107 are there in a 0.156 g sample of natural silver metal?

OpenStudy (anonymous):

\[0.156~g \times\frac{ 1~mol }{ 107~g }\times \frac{ 6.022 \times 10^{23}~atoms}{ 1~mol }\]

OpenStudy (littlebird):

I can't believe I didn't see this! :P

OpenStudy (littlebird):

But why doesn't isotope 109 affect the problem? Because I thought the problem implied that both isotopes are in the sample, and I have to figure out how much of the .156 g is isotope 107.

OpenStudy (anonymous):

I think isotope 109 is a detractor. The problem doesn't give any information about it, other than its existence. There's no way to include it in the calculation that I can see

OpenStudy (littlebird):

ok :)